# Bi-Hölder invariants in o-minimal structures

### v1

@@ 5. Homotopy types of the links of the tangent cones
$ST_{d_{3}}(X,C)\subset ST_{d_{1}}(Y,C)\subset\varphi(ST_{d}(X,C))\subset ST_{d_{2}}(Y,C).$
Note that $d_{3}>d_{1}>d_{2}$ and all tend to $d$ as $\alpha$ tends to $1$. Taking $\alpha_{0}$ close enough to $1$ such that for all $\alpha\in[\alpha_{0},1]$, $d_{i}\in(1,d_{0})$ for $i=1,2,3$. Consider the following commutative diagram induced by inclusion maps:
$\pi_{*}(ST_{d_{3}}(X,C)\setminus\{0\})$$\pi_{*}(ST_{d}(X,C)\setminus\{0\})$$\pi_{*}(ST_{d_{1}}(Y,C)\setminus\{0\})$$\pi_{*}(ST_{d_{2}}(Y,C)\setminus\{0\})$$\scriptstyle{\iota_{1*}}$$\scriptstyle{\iota_{2*}}$
Since $\iota_{i,*}$, $i=1,2$ are isomorphic, so are the others. This implies that the homotopy groups of $ST_{d}(X,C)\setminus\{0\}$ and $ST_{d^{\prime}}(Y,C)\setminus\{0\}$ are isomorphic for any $C>0$ and any $d,d^{\prime}\in(1,d_{0})$. It is known that $ST_{d}(*,C)\setminus\{0\}$, $\in\{X,Y\}$ can always deformation retracts onto ${\rm Lk}(ST_{d}(*,C),0)$. Therefore, ${\rm Lk}(ST_{d}(X,C),0)$ and ${\rm Lk}(ST_{d}(Y,C),0)$ are of the same homotopy type. The proof is complete. ∎
@@
Let $X\subset\mathbb{R}^{n}$ be a closed definable germ at 0. Then, there exists $d_{0}>1$ such that for every $C>0$ the homotopy types of ${\rm Lk}(ST_{d}(X,C),0)$ and ${\rm Lk}(ST_{d}(C_{0}(X),C),0)$ are the same for every $0<d<d_{0}$.
@@
Let $A:=C_{0}(X)$ and fix $C>0$. Define
$\Omega_{X}:=\{(d,x)\in(1,\infty)\times\mathbb{R}^{n}:{\rm dist}(x,X)\leq C\|x\|^{d}\},$
$\Omega_{A}:=\{(d,x)\in(1,\infty)\times\mathbb{R}^{n}:{\rm dist}(x,A)\leq C\|x\|^{d}\}.$
Consider the projection $\rho(d,x)=d$. By Hardt’s triviality theorem, there exists $d_{0}>1$ such that $\rho|_{\Omega_{\star}}$ is definably trivial over $(1,d_{0})$ for $\star\in\{X,A\}$.
Since $X$ and $A$ are $ST$-equivalent, we may choose
@@
The definable triviality of $\rho$ implies that there exists a deformation retract from $ST_{d_{2}}(X,C)$ onto $ST_{d_{4}}(X,C)$, and likewise a deformation retract from $ST_{d_{1}}(A,C)$ onto $ST_{d_{3}}(A,C)$. Moreover, these deformation retracts fix the origin. After removing the origin, and using the same inclusion argument as in the proof of Proposition 5.2, we conclude that the homotopy groups of $ST_{d}(X,C)\setminus\{0\}$ and $ST_{d}(A,C)\setminus\{0\}$ for every $1<d<d_{0}$ are isomorphic. This completes the proof. ∎
@@
Let $(X,0)$ and $(Y,0)$ be closed definable germs in $\mathbb{R}^{n}$. Then, there exists $\alpha_{0}\in(0,1)$ such that if $(X,0)$ and $(Y,0)$ are embedded bi-$\alpha$-Hölder equivalent for some $\alpha\in[\alpha_{0},1]$, then the homotopy types ${\rm Lk}(C_{0}(X),0)$ and ${\rm Lk}(C_{0}(Y),0)$ are the same.
@@
By Propositions 5.3 and 5.2, there are $0<\alpha_{0}<1$ and $d_{0}>1$ such that if $(X,0)$ and $(Y,0)$ are bi-$\alpha$-Hölder equivalent for some $\alpha\in[\alpha_{0},1]$ then the links of $ST_{d}(C_{0}(X),C)$ and $ST_{d}(C_{0}(Y),C)$ are of the same homotopy type for any $C>0$ and $d\in(0,d_{0})$. The result then follows by Lemma 5.1. ∎
@@
Let $(X,0)\subset\mathbb{R}^{n}$ and $(Y,0)\subset\mathbb{R}^{m}$ be closed definable germs. Then, there exists $\alpha_{0}\in(0,1)$ such that if $(X,0)$ and $(Y,0)$ are bi-$\alpha$-Hölder equivalent for some $\alpha\in[\alpha_{0},1]$, then the homotopy groups of ${\rm Lk}(C_{0}(X),0)$ and ${\rm Lk}(C_{0}(Y),0)$ are isomorphic.
@@ 5.1. Applications
Let $(X,0)\subset\mathbb{C}^{n}$ be complex analytic germ. If $(X,0)$ and $(\mathbb{R}^{k},0)$ are bi-$\alpha$-Hölder equivalent for every $0<\alpha<1$, then $X$ is smooth at 0.
@@
Let $(X,0)\subset\mathbb{C}^{n}$ be a complex analytic germ. Then, there exists $\alpha_{0}\in(0,1)$ such that if $(X,0)$ and $(\mathbb{R}^{k},0)$ are bi-$\alpha$-Hölder homeomorphic for some $\alpha\in[\alpha_{0},1]$, then $(X,0)$ is smooth.

### v2

@@ 5. Homotopy types of the links of the tangent cones
$\varphi(ST_{d_{3}}(X,C))\subset ST_{d_{1}}(Y,C)\subset\varphi(ST_{d}(X,C))\subset ST_{d_{2}}(Y,C).$
Note that $d_{3}>d_{1}>d_{2}$ all tend to $d$ as $\alpha$ tends to $1$. Take $\alpha_{0}$ close enough to $1$ such that for all $\alpha\in[\alpha_{0},1]$, $d_{i}\in(1,d_{0})$ for $i=1,2,3$. Consider the following commutative diagram induced by inclusion maps:
$\pi_{*}\bigg(\varphi(ST_{d_{3}}(X,C)\setminus\{0\})\bigg)$$\pi_{*}\bigg(\varphi(ST_{d}(X,C)\setminus\{0\})\bigg)$$\pi_{*}\bigg(ST_{d_{1}}(Y,C)\setminus\{0\}\bigg)$$\pi_{*}\bigg(ST_{d_{2}}(Y,C)\setminus\{0\}\bigg)$$\scriptstyle{\iota_{1*}}$$\scriptstyle{\iota_{2*}}$
Since $\iota_{i,*}$, $i=1,2$ are isomorphic, so are the others in the diagram. As $\varphi$ is a homeomorphism, it follows that the homotopy groups of $ST_{d_{3}}(X,C)\setminus\{0\}$ and $ST_{d_{1}}(Y,C)\setminus\{0\}$ are isomorphic. Thus, the homotopy types of $ST_{d^{\prime}}(X,C)\setminus\{0\}$ and $ST_{d^{\prime\prime}}(Y,C)\setminus\{0\}$ are the same for all $d^{\prime},d^{\prime\prime}\in(1,d_{0})$. It is known that $ST_{d}(Z,C)\setminus\{0\}$, $Z=X,Y$ can always deformation retracts onto ${\rm Lk}(ST_{d}(Z,C),0)$. Therefore, ${\rm Lk}(ST_{d}(X,C),0)$ and ${\rm Lk}(ST_{d}(Y,C),0)$ are of the same homotopy type.
Choose $\alpha_{0}=\alpha_{0}(d)$ for any $d$. We see that $\alpha_{0}$ satisfies the conclusion of the proposition. The proof is complete. ∎
@@
Let $X\subset\mathbb{R}^{n}$ be a closed definable germ at 0. For every $C>0$, there exists constant $d_{0}>1$ such that for every $0<d<d_{0}$ the homotopy types of ${\rm Lk}(ST_{d}(X,C),0)$ and ${\rm Lk}(ST_{d}(C_{0}(X),C),0)$ are the same.
@@
Let $A$ be the germ of $C_{0}(X)$ as 0 and fix $C>0$. Define
$\Omega_{X}:=\{x\in B(0,1):{\rm dist}(x,X)\leq C\|x\|^{d},d\in(1,2]\},$
$\Omega_{A}:=\{x\in B(0,1):{\rm dist}(x,A)\leq C\|x\|^{d},d\in(1,2]\}.$
Consider the continuous map $\rho_{X}=\sup\{d\in(1,2]:{\rm dist}(x,X)\leq C\|x\|^{d}\}$ and $\rho_{A}=\sup\{d\in(1,2]:{\rm dist}(x,A)\leq C\|x\|^{d}\}$ defined on $\Omega_{X}\setminus X$ and $\Omega_{A}\setminus A$ respectively. Arguing as in the proof of Proposition 5.2, we obtain a constant $d_{0}>1$ such that for any $1<d<d^{\prime}<d_{0}$ there are strong deformation retracts from $ST_{d}(X,C)\setminus\{0\}$ onto $ST_{d^{\prime}}(X,C)\setminus\{0\}$ and from $ST_{d}(A,C)\setminus\{0\}$ onto $ST_{d^{\prime}}(A,C)\setminus\{0\}$.
Since $X$ and $A$ are $ST$-equivalent, we may choose numbers
@@
After removing the origin, and using the same inclusion argument used in Proposition 5.2, we conclude that the homotopy groups of $ST_{d}(X,C)\setminus\{0\}$ and $ST_{d}(A,C)\setminus\{0\}$ for every $1<d<d_{0}$ are isomorphic. This completes the proof. ∎
@@
Let $(X,0)$ and $(Y,0)$ be closed definable germs in $\mathbb{R}^{n}$. Then, there exists $\alpha_{0}\in(0,1)$ such that if $(X,0)$ and $(Y,0)$ are embedded bi-$\alpha$-Hölder equivalent for some $\alpha\in[\alpha_{0},1]$, then the homotopy types of ${\rm Lk}(C_{0}(X),0)$ and ${\rm Lk}(C_{0}(Y),0)$ are the same.
@@
Fix a constant $C>0$. By Propositions 5.3 and 5.2, there exist $d_{0}>1$ and $\alpha_{0}\in(0,1)$ such that if $(X,0)$ and $(Y,0)$ are embedded bi-$\alpha$-Hölder equivalent for some $\alpha\in[\alpha_{0},1]$, then for every $d\in(1,d_{0})$ the links
${\rm Lk}\!\big(ST_{d}(C_{0}(X),C),0\big)\quad\text{and}\quad{\rm Lk}\!\big(ST_{d}(C_{0}(Y),C),0\big)$
have the same homotopy type.
By Lemma 5.1, the homotopy groups of ${\rm Lk}(ST_{d}(C_{0}(X),C),0)$ and ${\rm Lk}(C_{0}(X),0)$ are isomorphic, and likewise the homotopy groups of ${\rm Lk}(ST_{d}(C_{0}(Y),C),0)$ and ${\rm Lk}(C_{0}(Y),0)$ are isomorphic. It follows that ${\rm Lk}(C_{0}(X),0)$ and ${\rm Lk}(C_{0}(Y),0)$ have the same homotopy type.
Therefore, any such $\alpha_{0}$ satisfies the conclusion of the theorem. ∎
@@
Let $(X,0)\subset\mathbb{R}^{n}$ and $(Y,0)\subset\mathbb{R}^{m}$ be closed definable germs. Then, there exists $\alpha_{0}\in(0,1)$ such that if $(X,0)$ and $(Y,0)$ are bi-$\alpha$-Hölder equivalent for some $\alpha\in[\alpha_{0},1]$, then the homotopy type of ${\rm Lk}(C_{0}(X),0)$ and ${\rm Lk}(C_{0}(Y),0)$ are the same.
@@ 6. Applications
Let $(X,0)\subset\mathbb{C}^{n}$ be complex analytic germ. If $(X,0)$ and $(\mathbb{R}^{k},0)$ are bi-$\alpha$-Hölder equivalent for every $0<\alpha<1$, then $(X,0)$ is smooth.
@@
Let $(X,0)\subset\mathbb{C}^{n}$ be a complex analytic germ. Then, there exists $\alpha_{0}\in(0,1)$ such that if $(X,0)$ and $(\mathbb{R}^{k},0)$ are bi-$\alpha$-Hölder equivalent for some $\alpha\in[\alpha_{0},1]$, then $(X,0)$ is smooth.